# KnobGen: Controlling the Sophistication of Artwork in Sketch-Based Diffusion Models

### v2

@@ Abstract
Recent advances in diffusion models have significantly improved text-to-image (T2I) generation, but they often struggle to balance fine-grained precision with high-level control. Methods like ControlNet and T2I-Adapter excel at following sketches by seasoned artists but tend to be overly rigid, replicating unintentional flaws in sketches from novice users. Meanwhile, coarse-grained methods, such as sketch-based abstraction frameworks, offer more accessible input handling but lack the precise control needed for detailed, professional use. To address these limitations, we propose KnobGen, a dual-pathway framework that democratizes sketch-based image generation by seamlessly adapting to varying levels of sketch complexity and user skill. KnobGen uses a Coarse-Grained Controller (CGC) module for high-level semantics and a Fine-Grained Controller (FGC) module for detailed refinement. The relative strength of these two modules can be adjusted through our knob inference mechanism to align with the user’s specific needs. These mechanisms ensure that KnobGen can flexibly generate images from both novice sketches and those drawn by seasoned artists. This maintains control over the final output while preserving the natural appearance of the image, as evidenced on the MultiGen-20M dataset and a newly collected sketch dataset.
@@ 1 Introduction
Diffusion models (DMs) have revolutionized text-to-image (T2I) generation by generating visually rich images based on text prompts, excelling at capturing various levels of detail—from textures to high-level semantics (Saharia et al., 2022; Rombach et al., 2022; Nichol et al., 2021; Ramesh et al., 2021; Navard & Yilmaz, 2024). Despite their success, one of the primary limitations of these models is their inability to precisely convey spatial layout of the user-provided sketches. While text prompts can describe scenes, they struggle to capture complex spatial features, which makes it challenging to align generated images with user intent. This is particularly intensified when these users vary in skill and experience (Chowdhury et al., 2023; Song et al., 2017; Yang et al., 2023; Jiang et al., 2024).
To improve spatial control, sketch-conditioned DMs like ControlNet (Zhang et al., 2023), T2I-Adapter (Mou et al., 2024), and ControlNet++ (Li et al., 2024) have introduced mechanisms to allow users to input sketches that guide the generated image. However, these approaches primarily cater to artistic sketches with intricate details, which poses a challenge for novice users. When presented with rough sketches, these models rigidly align to unintentional flaws, producing results that misinterpret the user’s intent and fail to achieve the desired visual outcome. Furthermore, we observed that the quality and alignment of the generated images with the input sketch are highly sensitive to the weighting parameter that governs the model’s dependence on the condition, Figure 2.
@@
In contrast, some frameworks like (Koley et al., 2024) [^1] The code and model weights at the time of submission were unavailable, preventing result reproduction. have attempted to address the needs of novice users by introducing sketch abstraction. Although this democratizes the generation process,  Koley et al. (2024) is limited to covering only 125 categories of sketch subjects and cannot handle unseen categories, significantly limiting the generalizability of the pre-trained DM to a limited number of subjects. Moreover, its abstraction-aware framework is not suitable for artistic-level sketches whose purpose is to guide the DM to follow a particular spatial layout. Additionally, the removal of the text-based conditioning in DM makes these models ignore the semantic power provided by text in diffusion models trained on large-scale image-text pairs. Additionally, it limits their ability to differentiate between visually similar but semantically distinct objects- such as zebra and horse.
In a nutshell, existing methods for sketch-based image generation tend to focus on either end of the user-level spectrum. Professional-oriented models like ControlNet and T2I-Adapter are designed to handle only artistic-grade sketches Fig. 3.a, while amateur-oriented approaches Koley et al. (2024), cater to novice sketches without text guidance Fig. 3.b. These methods often fail to integrate both fine-grained and coarse-grained control, limiting their adaptability across different user types and sketch complexities.
@@
Figure 3: Comparison across various sketch-control in DM. (a) fine-grained control based method such as ControlNet or T2I-adapter rigidly resembles a novice sketch resulting in an unrealistic image (b) abstraction-aware frameworks such as Koley et al. (2024) fails to capture fine grained-detials without text guidance(c) while our proposed KnobGen smoothes out the imperfection of the user drawing and preserves the features of the novice sketch. FGC: Fine-grained Controller, CGC: Coarse-grained Controller, ET: Text Encoder, EI: Image Encoder, DM: Diffusion Model.
@@
To address these challenges, we propose KnobGen, a dual-pathway framework designed to empower a pre-trained DM with the capability to handle both professional and amateur-oriented approaches. KnobGen seamlessly integrates fine-grained and coarse-grained sketch control into a unified architecture, allowing it to adapt to varying levels of sketch complexity and user expertise. Our model is built on two key pathways, Macro Pathway and Micro Pathway. The Macro Pathway extracts the high-level visual and language semantics from the sketch image and the text prompt using CLIP encoders and injects them into the DM via our proposed Coarse-Grained Controller (CGC). The Micro Pathway injects low-level, detailed features and semantics directly from sketch through our Fine-Grained Controller (FGC) module.
@@
Additionally, we propose two new approaches for training and inference in order to maintain a robust control of the Micro and Macro Pathways in the conditional generation. First, we introduce Modulator, a mechanism dynamically adjusting the influence of the FGC during training, ensuring that the CGC dominates in the early training phase to prevent overfitting to low-level sketch features extracted by the FGC module. This allows the model to optimally rely on both Pathways to capture high- and low-level spatial and semantic features. At inference, the Knob mechanism offers user-driven control during denoising steps, allowing adjustment of the level of fidelity between the generated image and the user’s inputs- sketch and text- by manipulating Micro and Macro Pathways. These new training and inference approaches ensure that KnobGen effectively handles not only novice sketches but also artistic-grade ones, adapting to varying levels of sketch complexity and user preferences.
Our key contributions are as follows:
  • Dual-Pathway Framework for Sketch-Based Image Generation: KnobGen proposes an add-on framework to DMs that handles image generation flexibly for a broad spectrum of users, from novice to seasoned artists, through our proposed micro-macro path.
  • Training Modulator for Balanced Coarse-Fine Grained Incorporation: Our dynamic modulator regulates the influence of our CGC and FGC modules throughout the training. This prevents premature overfitting to fine details, allowing the model to first capture high-level spatial coherence before gradually introducing more specific features.
  • Inference Knob for Adaptive Image Generation: Our inference-time Knob mechanism allows users to control the level of fidelity between the generated image and the inputs, i.e. sketch image and text. Our Knob mechanism tweaks the level of abstraction and details of the generated image adaptively based on the user’s preference.
@@ 2.1 Diffusion Models
Recent advances in DM have enabled high-quality image generation with improved sample diversity (Ho et al., 2020; Dhariwal & Nichol, 2021; Ho & Salimans, 2021; Nichol et al., 2021; Saharia et al., 2022; Shamshiri et al., 2024; Ramesh et al., 2022; Perera et al., 2024; Peebles & Xie, 2023), often exceeding the performance of Generative Adversarial Networks (GAN) (Goodfellow et al., 2014; Karras et al., 2019; 2021; Sauer et al., 2022). DMs are built on the concept of diffusion processes, where data are progressively corrupted by noise over several timesteps. The models learn to reverse this process by iteratively denoising noisy samples, transforming pure noise back into the original data distribution. Several studies, such as DDIM (Song et al., 2021), DPM-solver (Lu et al., 2022), and Progressive Distillation (Salimans & Ho, 2022), have focused on accelerating DMs’ generation process through more efficient sampling methodologies. To address the high computational costs of training and sampling, recent research has successfully employed strategies to project the original data into a lower-dimensional manifold, with DMs being trained within this latent space. Representative methods include LSGM (Vahdat et al., 2021), LDM (Rombach et al., 2022), and DALLE-2 (Ramesh et al., 2022), all of which leverage this latent space approach to improve efficiency while maintaining high generation quality.
Figure 4: KnobGen vs. baseline on novice sketches. KnobGen handles amateurish sketches by injecting features from the Micro and Macro Pathways in a controlled manner. Dual pathway design ensures that the generated image is faithful to the spatial layout of the original input sketch and the image has a natural appearance. Baseline methods, however, exhibit difficulty in maintaining these desired properties in their generations.
@@ 2.2 Text-to-Image Diffusion
In addition to producing high-quality and diverse samples, DMs offer superior controllability, especially when guided by textual prompts (Rombach et al., 2022; Xue et al., 2023; Chen et al., 2024; Podell et al., 2024; Esser et al., 2024). Imagen (Saharia et al., 2022) employs a pretrained large language model (e.g., T5 (Raffel et al., 2020)) and a cascade architecture to achieve high-resolution, photorealistic image generation. LDM (Rombach et al., 2022), also known as Stable Diffusion (SD), performs the diffusion process in the latent space with textual information injected into the underlying UNet through a cross-attention mechanism, allowing for reduced computational complexity and improved generation fidelity. To further address challenges when handling complex text prompts with multiple objects and object-attribution bindings, RPG (Yang et al., 2024) proposed a training-free framework that harnesses the chain-of-thought reasoning capabilities of multimodal large language models (LLMs) to enhance the compositionality of T2I generation. Ranni (Feng et al., 2024) tackles this problem by introducing a semantic panel that serves as an intermediary between text prompts and images; an LLM is finetuned to generate semantic panels from text which are then embedded and injected into the DM for direct composition. Our proposed method aligns with the SD paradigm but diverges by incorporating a composite module that combines textual information with coarse-grained information from sketch inputs, thereby injecting more comprehensive high-level semantics into the diffusion model.
@@ 2.3 Conditional Diffusion with Semantic Maps
As textual prompts often lack the ability to convey detailed information, recent research has explored conditioning DMs on more complex or fine-grained semantic maps, such as sketches, depth maps, normal maps, etc. Works such as T2I-Adapter (Mou et al., 2024), ControlNet (Zhang et al., 2023), and SCEdit (Jiang et al., 2024), leverage pretrained T2I models but employ different mechanisms to interpret and integrate these detailed conditions into the diffusion process. UniControl (Qin et al., 2023) proposes a task-aware module to unify $N$ different conditions (i.e. $N=9$) in a single network, achieving promising multi-condition generation with significantly fewer model parameters compared to a multi-ControlNet approach. While Koley et al. (2024) attempts to democratize sketch-based diffusion models, their approach faces several significant limitations, as discussed in the Introduction section. In contrast, our dual-pathway method integrates both fine-grained and coarse-grained sketch conditions while maintaining the option for textual prompts. This design offers greater flexibility and control, accommodating users ranging from amateurs to professionals.
@@ 3.1 Preliminary
Diffusion models (Ho et al., 2020) define a generative process by gradually adding noise to input data $z_{0}$ through a Markovian forward diffusion process $q(z_{t}|z_{0})$. At each timestep $t$, noise is introduced into the data as follows:
@@
where $\epsilon$ is sampled from a standard Gaussian distribution, and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$, with $\alpha_{t}=1-\beta_{t}$ representing a differentiable function of the timestep $t$. The diffusion process gradually converts $z_{0}$ into pure Gaussian noise $z_{T}$ over time.
@@
$\mathcal{L}(\epsilon_{\theta})=\sum_{t=1}^{T}\mathbb{E}_{z_{0}\sim q(z_{0}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\|\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon)-\epsilon\|_{2}^{2}\right].$ (2)
In controllable generation tasks (Zhang et al., 2023; Mou et al., 2024), where both image condition $c_{v}$ and text prompt $c_{t}$ are provided, the diffusion loss function can be extended to include these conditioning inputs. The loss at timestep $t$ is modified as:
@@
During inference, given an initial noise vector $z_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the final image $x_{0}$ is recovered through a step-by-step denoising process (Ho et al., 2020), where the denoised estimate at each step $t$ is calculated as:
@@
with $\epsilon_{\theta}$ being the noise predicted by the U-Net (Ronneberger et al., 2015) at timestep $t$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ representing the variance of the posterior Gaussian distribution $p_{\theta}(z_{0})$. This iterative process gradually refines $z_{t}$ until it converges to the denoised image $z_{0}$.
@@ 3.2 Dual Pathway
Our model introduces a dual-pathway framework that harmonizes high-level semantic abstraction with precise, low-level control over visual details, Figure 5. The integration of the CGC module and FGC module enables KnobGen to adaptively inject high-level semantics and low-level features throughout the denoising process. This design ensures that the model can scale its output complexity based on user input, thus supporting a wide spectrum of sketch sophistication levels, from amateur to professional-grade sketches.

### v3

@@ Abstract
Recent advances in diffusion models have significantly improved text-to-image (T2I) generation, but they often struggle to balance fine-grained precision with high-level control. Methods like ControlNet and T2I-Adapter excel at following sketches by seasoned artists but tend to be replicating unintentional flaws in sketches from novice users. Meanwhile, coarse-grained methods, such as sketch-based abstraction frameworks, offer more accessible input handling but lack the precise control needed for professional use. To address these limitations, we propose KnobGen, a dual-pathway framework that democratizes sketch-based image generation by adapting to varying levels of sketch complexity and user skill. KnobGen uses a Coarse-Grained Controller (CGC) module for high-level semantics and a Fine-Grained Controller (FGC) module for detailed refinement. The relative strength of these two modules can be adjusted through our knob inference mechanism to align with the user’s specific needs. These mechanisms ensure that KnobGen can flexibly generate images from both novice sketches and those drawn by seasoned artists. This maintains control over the final output while preserving the natural appearance of the image, as evidenced on the MultiGen-20M dataset and a newly collected sketch dataset. https://github.com/aminK8/KnobGen
@@ 1 Introduction
Diffusion models (DMs) have revolutionized text-to-image (T2I) generation by generating visually rich images based on text prompts, excelling at capturing various levels of detail—from textures to high-level semantics [38, 36, 26, 34, 25]. Despite their success, one of the primary limitations of these models is their inability to precisely convey spatial layout of the user-provided sketches. While text prompts can describe scenes, they struggle to capture complex spatial features, which makes it challenging to align generated images with user intent. This is particularly intensified when these users vary in skill and experience [3, 43, 52, 13].
@@
To improve spatial control, sketch-conditioned DMs like ControlNet [54], T2I-Adapter [24], and ControlNet++ [19] have introduced mechanisms to allow users to input sketches that guide the generated image. However, these approaches primarily cater to artistic sketches with intricate details, which poses a challenge for novice users. When presented with rough sketches, these models rigidly align to unintentional flaws, producing results that misinterpret the user’s intent and fail to achieve the desired visual outcome. Furthermore, as shown in Figure 2 we observed that the quality and alignment of the generated images with the input sketch are highly sensitive to the weighting parameter that governs the model’s dependence on the condition.
@@
In contrast, some frameworks such as [18, 47] [^1] The codes and model weights at the time of submission were unavailable which prevents reproducibility. have attempted to address the needs of novice users by introducing sketch abstraction. Although this democratizes the generation process,  [18] is limited to covering only 125 categories of sketch subjects and cannot handle unseen categories, significantly limiting the generalizability of the pre-trained DM to a limited number of subjects. Moreover, its abstraction-aware framework is not suitable for artistic-level sketches whose purpose is to guide the DM to follow a particular spatial layout. Additionally, the removal of the text-based conditioning in DM makes these models ignore the semantic power provided by text in diffusion models trained on large-scale image-text pairs. Additionally, it limits their ability to differentiate between visually similar but semantically distinct objects- such as zebra and horse. In  [47] the latent space of DM is rigidly aligned with that of the sketch, resulting in maximal reliance on the input sketch.
@@
Figure 3: Comparison across various sketch-control in DM. (a) fine-grained control based method such as ControlNet or T2I-adapter rigidly resembles a novice sketch resulting in an unrealistic image (b) abstraction-aware frameworks such as [18] fails to capture fine grained-detials without text guidance(c) while our proposed KnobGen smoothes out the imperfection of the user drawing and preserves the features of the novice sketch. FGC: Fine-grained Controller, CGC: Coarse-grained Controller, ET: Text Encoder, EI: Image Encoder, DM: Diffusion Model.
In a nutshell, existing methods for sketch-based image generation tend to focus on either end of the user-level spectrum. As illustrated in Figure 3.a ,fine-grained conditioning modules like ControlNet and T2I-Adapter are designed to handle only artistic-grade sketches, while amateur-oriented approaches [18] in Figure 3.b cater to novice sketches without text guidance. These methods often fail to integrate both fine-grained and coarse-grained control across different user types and sketch complexities.
@@
To address these challenges, we propose KnobGen, a dual-pathway framework designed to empower a pre-trained DM with the capability to handle both professional and amateur-oriented approaches. KnobGen seamlessly integrates fine-grained and coarse-grained sketch control into a unified architecture, allowing it to adapt to varying levels of sketch complexity and user expertise. Our model is built on two key pathways, Macro Pathway and Micro Pathway. The Macro Pathway extracts the high-level visual and language semantics from the sketch image and the text prompt using CLIP encoders and injects them into the DM via our proposed Coarse-Grained Controller (CGC). The Micro Pathway injects low-level features directly from sketch through our Fine-Grained Controller (FGC).
@@
Additionally, we propose two new approaches for training and inference in order to maintain a robust control of the Micro and Macro Pathways in the conditional generation. First, we introduce Modulator, a mechanism dynamically adjusting the influence of the FGC during training, ensuring that the CGC dominates in the early training phase to prevent overfitting to low-level sketch features extracted by the FGC module. This allows the model to optimally rely on both Pathways to capture high- and low-level spatial and semantic features. At inference, the Knob mechanism offers user-driven control during denoising steps, allowing adjustment of the level of fidelity between the generated image and the user’s inputs- sketch and text- by manipulating Micro and Macro Pathways. These new training and inference approaches ensure that KnobGen effectively handles not only novice sketches but also artistic-grade ones. Our key contributions are as follows:
  • Dual-Pathway Framework for Adaptive Sketch-Based Image Generation: KnobGen introduces a novel dual-pathway design that balances fine-grained and coarse-grained pathways, providing controlled flexibility for sketches with varied levels of details. This integration extends KnobGen’s applicability across diverse user types, from novice sketchers to seasoned artists, addressing a major gap in prior sketch-guided DM.
  • Dynamic Modulator to Harmonize Coarse and Fine Detail During Training: Our modulator mechanism tunes the influence of coarse and fine-grained pathways throughout training, overcoming the tendency of fine-grained details to dominate early stages. By balancing these inputs, our approach achieves optimal spatial layout and feature refinement, which SOTA models lack due to their reliance on fixed weighting schemes.
  • Inference-Time Knob for User-Controlled Sketch Fidelity and Realism: Unlike uniform weighting schemes in other models, our knob mechanism variably introduces detail based on user input, preserving both spatial adherence to sketches and natural image appearance. This user-driven control allows KnobGen to flexibly adapt to diverse user needs.
@@ 2.1 Diffusion Models
Recent advances in DM have enabled high-quality image generation with improved sample diversity [11, 4, 10, 26, 38, 35, 29, 42, 14, 28, 23], often exceeding the performance of Generative Adversarial Networks (GAN) [7, 15, 16, 40]. DMs are built on the concept of diffusion processes, where data are progressively corrupted by noise over several timesteps. The models learn to reverse this process by iteratively denoising noisy samples, transforming pure noise back into the original data distribution. Several studies, such as DDIM [44], DPM-solver [20], and Progressive Distillation [39], have focused on accelerating DMs’ generation process through more efficient sampling methodologies. To address the high computational costs of training and sampling, recent research has successfully employed strategies to project the original data into a lower-dimensional manifold, with DMs being trained within this latent space. Representative methods include LSGM [46], LDM [36], and DALLE-2 [35] which leverage latent space.
@@ 2.2 Text-to-Image Diffusion
In addition to producing high-quality and diverse samples, DMs offer superior controllability, especially when guided by textual prompts [36, 50, 2, 30, 41, 5]. Imagen [38] employs a pretrained large language model (e.g., T5 [33]) and a cascade architecture to achieve high-resolution, photorealistic image generation. LDM [36], also known as Stable Diffusion (SD), performs the diffusion process in the latent space with textual information injected into the underlying UNet through a cross-attention mechanism, allowing for reduced computational complexity and improved generation fidelity. To further address challenges when handling complex text prompts with multiple objects and object-attribution bindings, RPG [53] proposed a training-free framework that harnesses the chain-of-thought reasoning capabilities of multimodal large language models (LLMs) to enhance the compositionality of T2I generation. Ranni [6] tackles this problem by introducing a semantic panel that serves as an intermediary between text prompts and images; an LLM is finetuned to generate semantic panels from text which are then embedded and injected into the DM for direct composition. Our proposed method aligns with the SD paradigm but diverges by incorporating a composite module that combines textual information with coarse-grained information from sketch inputs, thereby injecting more comprehensive high-level semantics into the diffusion model.
Figure 4: KnobGen vs. baseline on novice sketches. KnobGen handles novice sketches by injecting features from the Micro and Macro Pathways in a controlled manner. Dual pathway design ensures that the generated image is faithful to the spatial layout of the original input sketch and the image has a natural appearance. Baseline methods, however, exhibit difficulty in maintaining these desired properties in their generations. We also provide examples with null prompt as an ablation study to demonstrate the robustness of KnobGen.
@@ 2.3 Conditional Diffusion with Semantic Maps
As textual prompts often lack the ability to convey detailed information, recent research has explored conditioning DMs on more complex or fine-grained semantic maps, such as sketches, depth maps, normal maps, etc. Works such as T2I-Adapter [24], ControlNet [54], and SCEdit [13], leverage pretrained T2I models but employ different mechanisms to interpret and integrate these detailed conditions into the diffusion process. UniControl [31] proposes a task-aware module to unify $N$ different conditions (i.e. $N=9$) in a single network, achieving promising multi-condition generation with significantly fewer model parameters compared to a multi-ControlNet approach. While [18] attempts to democratize sketch-based diffusion models, their approach faces several significant limitations, as discussed in the Introduction section. In contrast, our dual-pathway method integrates both fine-grained and coarse-grained sketch conditions while maintaining the option for textual prompts.
@@ 3 Method
The design of KnobGen ensures that low- and high-level details from the conditional signal, i.e. the sketch, are incorporated in a balanced manner, both during training and inference. In section 3.3, we introduce the modulator, illustrated in Figure 5.A, which harmonizes the influence of fine-grained control in the training phase. The modulator prevents the fine-grained control from overpowering the coarse-grained control signal in the early training stages—a common challenge in generative models [51] that SOTA diffusion models often overlook. In section 3.4, we further describe our knob mechanism, Figure 5.C, which adaptively adjusts the level of detail during inference to align with the user’s skill level.
@@ 3.1 Preliminary
Diffusion models [11] define a generative process by gradually adding noise to input data $z_{0}$ through a Markovian forward diffusion process $q(z_{t}|z_{0})$. At each timestep $t$, noise is introduced into the data as follows:
@@
where $\epsilon$ is sampled from a standard Gaussian distribution, and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$, with $\alpha_{t}=1-\beta_{t}$ representing a differentiable function of the timestep $t$. The diffusion process converts $z_{0}$ into pure Gaussian noise $z_{T}$ over time.
@@
$\mathcal{L}(\epsilon_{\theta})=\sum_{t=1}^{T}\mathbb{E}_{z_{0}\sim q(z_{0}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\|\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}z_{0}\right.\\ \left.+\sqrt{1-\bar{\alpha}_{t}}\epsilon)-\epsilon\|_{2}^{2}\right].$ (2)
In controllable generation tasks [54, 24], where both image condition $c_{v}$ and text prompt $c_{t}$ are provided, the diffusion loss function can be extended to include these conditioning inputs. The loss at timestep $t$ is modified as:
@@
During inference, given an initial noise vector $z_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the final image $x_{0}$ is recovered through a step-by-step denoising process [11], where the denoised estimate at each step $t$ is calculated as:
@@
with $\epsilon_{\theta}$ being the noise predicted by the U-Net [37] at timestep $t$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ representing the variance of the posterior Gaussian distribution $p_{\theta}(z_{0})$. This iterative process gradually refines $z_{t}$ until it converges to the denoised image $z_{0}$.
@@ 3.2 Dual Pathway
Figure 5 demonstrates our model, a dual-pathway framework that harmonizes high-level semantic abstraction with precise, low-level control over visual details. The integration of the CGC module and FGC module enables KnobGen to adaptively inject high-level semantics and low-level features throughout the denoising process. This design ensures that the model can scale its output complexity based on user input, thus supporting a wide spectrum of sketch sophistication levels.